OpenStudy (anonymous):

What equations are parallel and perpendicular to y = -3x - 1 and that pass through the point (3, 1)?

OpenStudy (anonymous):

m1=-3 ......... m2=-3 cause it is parallel since m=y-y1/x-x1 therefore -3=y-1/x-3 therefore -3x+9=y-1......so....... y=-3x+10 m1=-3.............m2=1/3 cause it is orthogonal since m=y-y1/x-x1 therefore 1/3=y-1/x-3 therefore 3y-3=x-1.......so.....3y=x+2